You are considering buying a new home. You will need to borrow 350,000 to purchase the home. A mortgage company offers you a 20-year fixed rate at 9% (as an APR with monthly compounding). If you borrow the money from this mortgage company, your monthly payment will be closest to: A. 18,910.72 B. 3,149.04 C. 38,341.27 D. 3,070.42 E, 2,635.35 F. None of the above is correct.

hi how do i solve this please explain
Answer #1

Loan amount = 350,000

Term of loan = 20 years

Number of monthly payments = n = 20 years * 12 = 240

Annual interest rate = 9%

Monthly interest rate =r = 9%/12 = 0.75% = 0.0075

Loan payment = (Loan amount*r) / {1-(1+r)-n}

Monthly loan payment = (350000*0.0075)/(1-1.0075-240)

Monthly loan payment = 2625/0.8336 = 3149.04

Hence, correct answer is B. 3149.04